OpenStudy (jamanorthi):

how do i write 3.148% as a fraction in simplest form and as a decimal?

OpenStudy (akashdeepdeb):

3.148% = 3.148 PER CENT 3.148/100 As a fraction = 3148/100000 [And then simplify it] As a decimal when you divide it by 100 it shifts 2 places to the left. 3.148/100 = 0.03148 Understood? :)

OpenStudy (jamanorthi):

how u get 3148/100000 as a fraction

OpenStudy (akashdeepdeb):

Multiply the numerator by 100 and also the denominator to get rid of the decimal point in the numerator and get a fraction. Understood? :)

OpenStudy (jamanorthi):

but if u do that then u will get 314800/10000